# 3D-ADAM: A Dataset for 3D Anomaly Detection in Advanced Manufacturing

### v1

@@ Abstract
Surface defects are one of the largest contributors to low yield in the manufacturing sector. Accurate and reliable detection of defects during the manufacturing process is therefore of great value across the sector. State-of-the-art approaches to automated defect detection yield impressive performance on current datasets, yet still fall short in real-world manufacturing settings and developing improved methods relies on large datasets representative of real-world scenarios. Unfortunately, high-quality, high-precision RGB+3D industrial anomaly detection datasets are scarce, and typically do not reflect real-world industrial deployment scenarios. To address this, we introduce 3D-ADAM (3D Anomaly Detection in Advanced Manufacturing), the first large-scale industry-relevant dataset for high-precision 3D Anomaly Detection. 3D-ADAM comprises 14,120 high-resolution scans across 217 unique parts, captured using 4 industrial depth imaging sensors. It includes 27,346 annotated defect instances from 12 categories, covering the breadth of industrial surface defects. 3D-ADAM uniquely captures an additional 8,110 annotations of machine element features, spanning the range of relevant mechanical design form factors. Unlike existing datasets, 3D-ADAM is captured in a real industrial environment with variations in part position and orientation, camera positioning, ambient lighting conditions, as well as partial occlusions. Our evaluation of SOTA models across various RGB+3D anomaly detection tasks demonstrates the significant challenge this dataset presents to current approaches. We further validated the industrial relevance and quality of the dataset through an expert labelling survey conducted by industry partners. By providing this challenging benchmark, 3D-ADAM aims to accelerate the development of robust 3D Anomaly Detection models capable of meeting the demands of modern manufacturing environments. We provide our dataset for accessibility at: https://huggingface.co/datasets/pmchard/3D-ADAM and benchmark source code at: https://github.com/PaulMcHard/3D-ADAMBench
1 Introduction
@@
The manufacturing sector constantly strives to reduce costs and improve quality and consistency through the use of advancements in technology. A key goal is to automatically control and monitor the entire manufacturing life-cycle of a given part. This includes detecting any anomalies or defects that may arise during the manufacturing process, and taking appropriate corrective action scheer2015 .
@@
Accurately and reliably detecting surface defects during the manufacturing process remains a significant challenge across the manufacturing industry apqc2024 , affecting over 42% of manufacturers. 2D Anomaly Detection methods are a mature and well established technology, with a diverse range of high-volume 2D Anomaly Detection Datasets available mvtec2021 SMAP2018 and a large body of work focussed on tasks such as few shot anomaly detection WinCLIP2023 AnomalyCLIP2023 and unsupervised anomaly detection DFM2024 ContextFlow++2024 . In contrast, 3D Anomaly Detection methods lack the same degree of maturity. We attribute this in part to the fact that suitable datasets are limited and do not provide the breadth of features and defect examples required for developing solutions capable of performing at an acceptable level in advanced manufacturing contexts. Three leading datasets have been proposed previously for the challenge of 3D Anomaly Detection, MVTec3D-AD mvtec2022 , Eyecandies eyecandies2022 and Real3D-AD real3dad2023 . Each of these datasets has several limitations which make them unsuitable for industrial applications, with the common issue between them being a lack of breadth in the defect and feature categories presented in manufacturing samples.
Hence, we introduce 3D-ADAM, a large-scale, high-resolution, multi-camera dataset for 3D-Anomaly Detection in Advanced Manufacturing. 3D-ADAM addresses the limitations of existing datasets by including comprehensive suite of examples of defects in 12 classes across a range of machine element features Mott2023 , which cover the range of surface defect types that may be encountered in precision part manufacturing environments. 3D-ADAM proposes a set of 16 machine elements Mott2023 features across all part categories, with care taken in curating a range of defect examples displayed on each machine element type.
The manufactured parts contained in 3D-ADAM constitute a complete assembly for a robot arm mechanism, based on the open-source BCN3D-Moveo system BCNMoveo2022 and augmented with additional components sourced from the Thingi10K Thingi10K2016 part dataset in order to ensure all relevant machine elements are adequately represented. We manufactured all the parts, both good and defective, in Polylactic Acid (PLA) as it is amongst the most commonly used materials for Fused Deposition Modelling (FDM) based additive manufacturing. Some defects naturally arose during the printing process, while others were forced by adjusting the printing parameters or by manual intervention. We assess the performance of our dataset in the tasks of 3D and RGB + 3D Anomaly Detection in a supervised setting across a suite of leading models, demonstrating that the 3D-ADAM dataset presents a novel challenge for current leading architectures in 3D and RGB+3D Anomaly Detection, which will bridge the gap between the current SOTA in 3D Anomaly Detection models and the level of performance required for deployment in real-world scenarios. Our main contributions are:
  • We propose the first multi-camera, multi-modal, industry-relevant dataset for 3D and RGB+3D Anomaly Detection and Localization. Comprising 14,120 high-resolution image and 3D point cloud scans across 28 categories, with 27,346 annotated defect regions in 12 classes and a further 27,346 annotated machine element features in 16 classes (Figure 1).
  • We evaluate leading methods in supervised 3D and RGB+3D Anomaly Detection and Localisation. Our initial benchmark demonstrates that these methods under-perform on our dataset, and that there is significant room for improvement for such methods to be suitable for real-world environments.
  • We evaluate the quality of our defect segmentation mask annotations through a survey of industry experts which shows that individuals with expert knowledge of surface defects in advanced manufacturing contexts strongly agree with the image-level quality and pixel-level accuracy of the annotations proposed in our dataset.
@@
2 Related Work
3D Anomaly Detection Datasets.
Three datasets have previously been proposed for 3D Anomaly Detection, MVTec3D-AD mvtec2022 , Eyecandies eyecandies2022 and Real3D-AD real3dad2023 . Of these, MVTec 3D-AD and Real3D-AD represent the current primary datasets used as benchmarks for the problem of 3D Anomaly Detection in manufacturing. The Eyecandies dataset eyecandies2022 , while a novel proposal in its own right, comprises a set of synthetically generated parts and defects on object instances that are unsuitable for industrial contexts, and is therefore not considered a viable dataset for such real-world applications. In their proposal for the Real 3D-AD Dataset, Liu et al. real3dad2023 claim their dataset is more suitable for achieving high-precision 3D Anomaly Detection. Indeed, this claim is supported by their notably higher point resolution and precision when compared to the MVTec 3D-AD dataset, which is a result of the difference in point-cloud acquisition hardware used. However, the MVTec dataset remains far more popular among new techniques published for 3D Anomaly Detection. Both datasets feature a similar number of object categories, 10 in the MVTec 3D-AD Dataset and 11 in the Real3D-AD Dataset, respectively, with the volume of pointcloud scan data per object category also comparable. The key distinction between the two is that the MVTec dataset presents a superior and far more rigorous ground truth annotation methodology and ontology, and for this reason remains the more widely adopted.
Industrial Anomaly Detection.
Advances in computer vision and deep learning are leading the technological drive in improvements to defect detection in the manufacturing industry henriksson2020 . Saeed et al saeed2021 , for example, demonstrate that frameworks such as Faster Regional-Convolutional Neural Networks ren2017 , Faster R-CNN, perform strongly in solving the task of fault detection and small object detection in an industrial environment. The foundational work in Deep Learning on 3D Point Clouds is the PointNet qi2017 architecture, which proposed a novel architecture for classification and segmentation tasks, taking sets of unordered 3D points from point clouds as input. Since PointNet was proposed, it has provided the basis from which many deep learning challenges have been handled in 3D point cloud data. Its successor network, PointNet++ qi2017plus , introduced the idea of applying PointNet recursively to nested partitions of the input point set. The introduction of this approach allows the network to learn local features of the part by exploiting space metrics to develop an understanding of the context surrounding a region. Not only does this underpin some of the state-of-the-art 3D Anomaly detection models shapeGuided2023 , but it has also served as the foundation for a number of novel techniques and architectures that look to expand beyond it.
The current state-of-the-art in 3D Anomaly Detection cpmf2024 3dsr2024 shapeGuided2023 transfusion2024 presents a number of highly capable models for unsupervised 3D anomaly detection. These models have all been trained, tested and validated on the MVTec 3D-AD Dataset mvtec2022 . Each of these models is underpinned by backbones built on the PointNet qi2017 and ResNet resnet2015 models, however vary significantly in the architectures proposed. Chu et al. shapeGuided2023 propose a model for Shape-Guided expert learning, in which two separate architectures are trained for 3D feature extraction, based on PointNet part segmentation, and 2D anomaly detection, based on ResNet-based resnet2015 RGB feature mapping, respectively. The 3D features then act as a guide for the RGB features to build dual memory banks from which inference can be conducted for anomaly detection in an unseen part. Zvartanik et al. 3dsr2024 propose the Depth-Aware Discrete Auto-Encoder (DADA) architecture, which takes both the RGB and Depth image data jointly as a single input, enabling the model to learn in a general discrete latent space. The DADA architecture is the core of their 3D Dual Subspace Reprojection Network (3DSR) 3dsr2024 , which is an extension of the authors’ previous work zavrtanik2022 , in using dual-encoder models for 2D anomaly detection. Conversely, a leading model in 3D Anomaly Detection proposed by Cao et al. cpmf2024 does not rely on raw RGB information as input in any capacity. Instead, they propose a Pseudo Multi-Modal Feature, in which 2D features are extracted by creating projections of the 3D point cloud data into 2D at a given number of view-points, and relying on pre-trained 2D image networks. Their model, based on ResNet, is similar to Chu et al.’s shapeGuided2023 Shape-Guided architecture. These multi-view features are then aggregated together with hand-crafted 3D features, to construct the memory banks from which inference can be performed for anomaly detection. Lastly, Fucka et al. transfusion2024 propose a model that leverages modern diffusion-based techniques in a transparency-based diffusion model for 3D Anomaly Detection, with promising results.
On the MVTec 3D-AD Dataset upon which the leading models are all trained, Fucka et al.’s TransFusion model transfusion2024 , currently has the hightest performing results on both Detection and Segmentation tasks in RGB+3D Anomaly Detection, whilst Chu et al.’s shapeGuided2023 Shape-Guided model performs best on Segmentation benchmarks, and Cao et al.’s 3dsr2024 CPMF model provides the strongest performance on Detection Benchmarks, amongst available models. It should be noted that this description excludes the Global Local Feature Matching (GLFM) model proposed by Cheng et al. cheng2025 , as while this model outperforms both Shape-Guided and CPMF at the time of writing, it is not yet publicly available.
3 The 3D-ADAM Dataset
The 3D-ADAM dataset consists of 12,192 scans distributed equally between 4 high-resolution industrial 3D sensors. The sensors deployed are MechMind LSR-L, MechMind Nano, Stereolabs Zed 2i, and Intel Realsense (as shown in Figure 2). The details of these sensors are described in Table 1. To obtain a comprehensive coverage of the technologies deployed in industrial vision applications, these sensors vary in quality, cost, capture resolution, and depth-sensing technology.
@@ 3 The 3D-ADAM Dataset
Table 1: Specifications of sensors used to capture the 3D-ADAM dataset.
@@
We aim to provide exhaustive coverage of the range of relevant defect and machine element types, such that we offer a diverse range of defect examples for industrial anomaly detection applications. To achieve this, 3D-ADAM contains scans of parts from 29 distinct object categories. The majority of these parts represent the complete assembly for a robot arm mechanism, obtained from the open-source BCN3D-Moveo system BCNMoveo2022 . This set is augmented with additional components sourced from the open-source Thingi10k dataset Thingi10K2016 to provide a full suite of machine elements, covering common elements such as faces and edges, internal and external fillets, internal and external chamfers, holes, kerfs, tapers, indents, counterbores, countersinks as well as the most common gear types; spur gear teeth, rack gear teeth, spiral gear teeth and both clockwise and counter-clockwise helical gear teeth, for a total of 16 distinct machine element classes (as shown in Figure 1 in the supplementary material). In addition, we curate a full suite of defect examples across this breadth of machine element examples, distributed across the range of part categories. The range of defect classes is designed to represent the full spectrum of distinct surface defect types which can be encountered in manufacturing environments and includes (shown in Figure 1); cuts, bulges, holes, gaps, burrs, cracks, scratches, marks, warping, roughness and, specific to the additive manufacturing process employed in producing this dataset, over extrusion and under extrusion defects, for a total of 12 defect classes, which is the largest range provided by any dataset to date. Table 2 describes the breakdown of total parts, total scans and total number of defect and machine element annotations, as well as the number of unique defect classes and machine element classes featured in each category of our dataset.
Category Parts Images Defects Machine Elements # of Defect Classes # of Machine Element Classes
@@
Table 2: Statistical Breakdown. # of defect and # of machine element classes is given per per part category.
3.1 Structure & Data Acquisition
@@
We employed additive manufacturing via Fused Deposition Modelling (FDM) cano2021 to manufacture the parts included in our dataset. For this task, four distinct systems were employed: two Prusa i3 systems, an Ender 3 Pro, and an Ultimaker 2, and all parts were manufactured in Polylactic Acid (PLA). In the majority of instances, these systems were used to produce one complete set of "good" parts, manufactured at optimal settings. Many of these "good" examples were ultimately found to exhibit some minor degree of surface defects, discovered during the annotation stage. The set of parts for defective cases was distributed randomly and evenly amongst the four systems. Defect instances were created either by manipulating manufacturing settings on the systems, in order to create manufacturing process-based defects such as bulges, gaps, warping, over extrusion and under extrusion. Surface defect examples of a mechanical nature, in contrast to those generated by process error, were generated by replicating the handling and manipulation errors. These examples result in a real manufacturing environment, including the generation of cuts, holes, cracks, marks, and scratches. The majority of instances of burrs and roughness defects were as a result of naturally occurring defective features on parts.
@@
Figure 2: Experimental setup in laboratory conditions, showing mouting of the MechMind Nano, Intel RealSense and Stereolabs Zed2i sensors on an ABB IRB 120 robot arm, oriented towards the mounting fixture. Not pictured: MechMind LSR-L camera, mounted out of view above the cell frame.
@@
Data was acquired in a real-world environment in the robotics applications development cell at HAL Robotics. (Figure 2). Within this cell, an ABB IRB 120 robot arm was used to mount and orient the MechMind-Nano, Intel RealSense and Stereolabs Zed2i sensors, while MechMind LSR-L sensor was fixed to the upper exterior of the cell enclosure. The positions of all four industrial 3D imaging sensors were static throughout the data acquisition process. This mounting configuration was used to ensure that all four sensors would be operating within their optimal scanning ranges, described in Table 1, oriented towards a mounting fixture that held the parts in place during scanning. The mounting fixture allowed 2-axis rotation, with a 360-degree rotation around its primary axis and a 90-degree range of motion in its secondary axis. In each scanning procedure, the part was fixed at the centre of this fixture, and allowed to rotate through a range of orientations, with a minimum of a full rotation of the primary axis at 45-degree intervals, with the secondary axis fixed at both 0 and 30 degrees, respectively. This provided, in most instances, a complete view of the part. In those instances where additional coverage was desired, further variance in the secondary axis position was used to achieve the complete view.
Each of the industrial scanners deployed captured six-channel images in the capture resolutions defined in Table 1, with these channels representing the r, g, b, x, y and z coordinates with respect to the sensors’ local reference coordinate frame. These [x, y, z] vectors represent the point cloud scan obtained by the sensor, and in all instances are captured as .PLY files. Additionally, for all sensors, the corresponding [r, g, b] values for each pixel are captured as a PNG RGB image, which enables us to have a 1-to-1 mapping between the RGB and XYZ image data for each scan. The scene captured within our environment was open to indirect natural lighting sources during the period of scanning, as well as diffuse indirect artificial light sources when natural illumination levels were insufficient for proper data acquisition conditions. Thus, as scanning was conducted throughout the day, across weather and seasonal change variations, the data captured represents a wide range of natural and artificial lighting conditions as can be observed in Figure 1 and 3.
@@
3.2 Ground Truth Annotations

### v2

# 3D-ADAM: A Dataset for 3D Anomaly Detection in Additive Manufacturing

@@ Abstract
Surface defects are a primary source of yield loss in manufacturing, yet existing anomaly detection methods often fail in real-world deployment due to limited and unrepresentative datasets. To overcome this, we introduce 3D-ADAM, a 3D Anomaly Detection in Additive Manufacturing dataset, that is the first large-scale, industry-relevant dataset for RGB+3D surface defect detection in additive manufacturing. 3D-ADAM comprises 14,120 high-resolution scans of 217 unique parts, captured with four industrial depth sensors, and includes 27,346 annotated defects across 12 categories along with 27,346 annotations of machine element features in 16 classes. 3D-ADAM is captured in a real industrial environment and as such reflects real production conditions, including variations in part placement, sensor positioning, lighting, and partial occlusion. Benchmarking state-of-the-art models demonstrates that 3D-ADAM presents substantial challenges beyond existing datasets. Validation through expert labelling surveys with industry partners further confirms its industrial relevance. By providing this benchmark, 3D-ADAM establishes a foundation for advancing robust 3D anomaly detection capable of meeting manufacturing demands.
I Introduction
@@
The manufacturing sector constantly strives to reduce costs and improve quality and consistency through the use of advancements in technology. A key goal is to automatically control and monitor the entire manufacturing life-cycle of a given part. This includes detecting any anomalies or defects that may arise during the manufacturing process, and taking appropriate corrective action [1].
@@
Accurately and reliably detecting surface defects during the manufacturing process remains a significant challenge across the manufacturing industry [2], affecting over 42% of manufacturers. 2D Anomaly Detection methods are a mature and well established technology, with a diverse range of high-volume 2D Anomaly Detection Datasets available [3] [4] and a large body of work focused on tasks such as few shot anomaly detection [5] [6] and unsupervised anomaly detection [7] [8]. In contrast, 3D Anomaly Detection methods lack the same degree of maturity. We attribute this to the fact that suitable datasets are limited and do not provide the breadth of features and defect examples required for developing solutions capable of performing at an acceptable level across the breadth of manufacturing contexts. Four leading datasets have been proposed previously for the challenge of 3D Anomaly Detection, MVTec3D-AD [11], Eyecandies [12], Real3D-AD [13] and PAD [14]. Each of these datasets has several limitations which make them unsuitable for industrial applications, with the common issue between them being a lack of breadth in the defect and feature categories presented in the manufacturing samples.
@@
Hence, we introduce 3D-ADAM, a large-scale, high-resolution, multi-camera dataset for 3D-Anomaly Detection in Additive Manufacturing. 3D-ADAM addresses the limitations of existing datasets by including comprehensive suite of examples of defects in 12 classes across a range of machine element features [15], which cover the range of surface defect types that may be encountered in additive manufacturing environments. 3D-ADAM proposes a set of 16 machine element features [15] across all part categories, with care taken in curating a range of defect examples displayed on each machine element type.
The manufactured parts contained in 3D-ADAM constitute a complete assembly for a robot arm mechanism, based on the open-source BCN3D-Moveo system [16] and augmented with additional components sourced from the Thingi10K [17] part dataset in order to ensure all relevant machine elements are adequately represented. We manufactured all the parts, both good and defective, in Polylactic Acid (PLA) as it is amongst the most commonly used materials for Fused Deposition Modelling (FDM) based additive manufacturing. Some defects naturally arose during the printing process, while others were forced by adjusting the printing parameters or by manual intervention. We assess the performance of our dataset in the tasks of 2D, 3D and RGB + 3D Anomaly Detection in an unsupervised setting across a suite of leading models, demonstrating that the 3D-ADAM dataset presents a novel challenge for current leading architectures in 3D and RGB+3D Anomaly Detection, which will bridge the gap between the current SOTA in 3D Anomaly Detection models and the level of performance required for deployment in real-world scenarios. Our main contributions are:
  • We propose the first multi-camera, multi-modal, industry-relevant dataset for 3D and RGB+3D Anomaly Detection and Segmentation. Comprising 14,120 high-resolution image and 3D point cloud scans across 28 categories, with 27,346 annotated defect regions in 12 classes and a further 27,346 annotated machine element features in 16 classes.
  • We evaluate foundational and leading methods in unsupervised 2D, 3D and RGB+3D Anomaly Detection and Segmentation tasks. Our initial benchmark demonstrates that these methods under-perform on our dataset, and that there is significant room for improvement for such methods to be suitable for real-world environments.
  • We evaluate the quality of our ground-truth defect annotations through a survey of industry experts which shows that individuals with expert knowledge of surface defects in additive manufacturing contexts strongly agree with the image-level quality and pixel-level accuracy of the annotations proposed in our dataset.
II Related Work
TABLE I: Specifications of sensors used to capture the 3D-ADAM dataset.
@@ II Related Work
II-A Additive Manufacturing Defects
Additive Manufacturing encompasses a breadth of fundamentally different processes. In their extensive review of additive manufacturing processes and their effects on part defects, de Pastre et al. [9] categorise these into seven distinct processes; Binder Jetting, Directed Energy Deposition, Material Extrusion, Material Jetting, Powder Bed Fusion, Sheet Lamination and Vat Photo-Polymerisation. In the context of industrial robotics applications, it is only Material Extrusion processes, in particular Fused Deposition Modelling (FDM), which are relevant. Within the bounds of FDM processes, literature concludes that geometry is of more significant consideration than material when considering both the likelihood and form of defects occurring during the manufacturing process [9], [10].
II-B 3D Anomaly Detection Datasets
Four datasets have previously been proposed for 3D Anomaly Detection, MVTec3D-AD [11], Eyecandies [12], Real3D-AD [13] and PAD [14]. Of these, MVTec 3D-AD and Real3D-AD represent the current leading datasets used as benchmarks for the problem of 3D Anomaly Detection in manufacturing. The Eyecandies dataset [12] is comprised of synthetically generated parts and defects on object instances that are unsuitable for industrial contexts, and is thus not viable for such real-world applications. Similarly, the PAD dataset [14], presents novel approaches to dataset acquisition for 3D Anomaly Detection, however the foundation of their dataset is assembled LEGO models, with defects relevant only to that context, with little extension into the broader manufacturing landscape. In the Real 3D-AD Dataset, Liu et al. [13] claim their dataset is more suitable for achieving high-precision 3D Anomaly Detection, supported by their notably higher point resolution and precision when compared to the MVTec 3D-AD dataset. However, the MVTec dataset remains the most popular among new techniques published for 3D Anomaly Detection. Both datasets feature a similar number of object categories, 10 in the MVTec 3D-AD Dataset and 11 in the Real3D-AD Dataset, respectively, with the volume of scan data per object category also comparable. The key distinction between the two is that the MVTec dataset presents a superior and more rigorous ground truth annotation methodology, and thus remains more widely adopted. Table II compares the size and complexity between our dataset, MVTec3D-AD and Real3D-AD.
TABLE II: Dataset Comparison between SOTA and 3D-ADAM
Dataset
MVTec3D-AD Real3D-AD 3D-ADAM (Ours)
# of Total Scans
4,147 1,254 14,120
# of Categories
10 12 28
# of Defect Classes
8 2 12
# of Defect Annotations
1,148 602 27,346
# of Machine Element
Classes
- - 16
# of Machine Element
Annotations
- - 27,346
Setting
Single-Sensor,
Multi-view
Single-Sensor,
Multi-view
Multi-Sensor,
Multi-view
# Sensors
1 1 4
Annotation Type
Single Mask Single Mask Per-Defect Mask
II-C Industrial Anomaly Detection
The current state-of-the-art in 3D Anomaly Detection [27] [29] [28] presents a number of highly capable models for unsupervised 3D anomaly detection. These models have all been trained, tested and validated on the MVTec 3D-AD Dataset [11]. Each of these models is underpinned by backbones built on the PointNet [22] and ResNet [25] models, however vary significantly in the architectures proposed.
Zvartanik et al. [27] propose the Depth-Aware Discrete Auto-Encoder (DADA) architecture, which takes both the RGB and Depth image data jointly as a single input, enabling the model to learn in a general discrete latent space. The DADA architecture is the core of their 3D Dual Subspace Reprojection Network (3DSR) [27], which is an extension of the authors’ previous work [31], in using dual-encoder models for 2D anomaly detection. Conversely, a leading model in 3D Anomaly Detection proposed by Cao et al. [26] does not rely on raw RGB information as input in any capacity. Instead, they propose a Pseudo Multi-Modal Feature, in which 2D features are extracted by creating projections of the 3D point cloud data into 2D at a number of view-points, relying on pre-trained 2D image networks. These multi-view features are then aggregated together with hand-crafted 3D features to construct the memory banks from which inference can be performed for anomaly detection. Fucka et al. [29] propose a model that leverages modern diffusion-based techniques in a transparency-based diffusion model for 3D Anomaly Detection. Lastly, Cheng et al. [28] have released the most recent and most promising model to date, which uses a PatchCore [30] based backbone to produce both global and local feature mappings, which their Adapted Point Transformer then utilises in conjunction with one another to perform Anomaly Detection and Segmentation predictions.
On the MVTec 3D-AD Dataset upon which the leading models are all trained, Cheng et al.’s GLFM model [28], currently has the hightest performing results on both Detection and Segmentation tasks in RGB+3D Anomaly Detection, whilst Fucka et al.’s [29] TransFusion model performs best on 3D Segmentation benchmarks, and Cao et al.’s [27] CPMF model provides the strongest performance on 3D Detection Benchmarks, amongst available models.
III The 3D-ADAM Dataset
The 3D-ADAM dataset consists of 14,120 scans distributed equally between 4 high-resolution industrial 3D sensors. The sensors deployed are MechMind LSR-L, MechMind Nano, Stereolabs Zed 2i, and Intel Realsense (as shown in Figure 2). The details of these sensors are described in Table I. To obtain a comprehensive coverage of the technologies deployed in industrial vision applications, these sensors vary in quality, cost, capture resolution, and depth-sensing technology.
@@ III The 3D-ADAM Dataset
We aim to provide exhaustive coverage of the range of relevant defect and machine element types, such that we offer a diverse range of defect examples for industrial anomaly detection applications. To achieve this, 3D-ADAM contains scans of parts from 28 distinct objects. The majority of these parts represent the complete assembly for a robot arm mechanism, obtained from the open-source BCN3D-Moveo system [16]. This set is augmented with additional components sourced from the open-source Thingi10k dataset [17] to provide a full suite of machine elements, covering common elements such as faces and edges, internal and external fillets, internal and external chamfers, holes, kerfs, tapers, indents, counterbores, countersinks as well as the most common gear types; spur gear teeth, rack gear teeth, spiral gear teeth and both clockwise and counter-clockwise helical gear teeth, for a total of 16 distinct machine element classes. Examples of these machine element types can be found in the accompanying video.
In addition, we curate a full suite of defect examples across this breadth of machine element examples, distributed across the range of part categories. The range of defect classes is designed to represent the full spectrum of distinct surface defect types which can be encountered in additive manufacturing environments [9], [10] and includes (shown in Figure 1); cuts, bulges, holes, gaps, burrs, cracks, scratches, marks, warping, roughness and, specific to the additive manufacturing process employed in producing this dataset, over extrusion and under extrusion defects, for a total of 12 defect classes, which is the largest range provided by any dataset to date. Table III describes the breakdown of total parts, total scans and total number of defect and machine element annotations, as well as the number of unique defect classes and machine element classes featured in each category of our dataset.
TABLE III: Statistical Breakdown. # of defect and # of machine element classes is given per per part category.
Category Parts Images Defects
Machine
Elements
Defect
Classes
Machine Elem.
Classes
@@
III-A Structure & Data Acquisition
@@
We employed additive manufacturing via Fused Deposition Modelling (FDM) [18] to manufacture the parts included in our dataset. For this task, four distinct systems were employed: two Prusa i3 systems, an Ender 3 Pro, and an Ultimaker 2, and all parts were manufactured in Polylactic Acid (PLA). These systems were used to produce one complete set of ”good” parts, manufactured at optimal settings. Many of these ”good” examples were ultimately found to exhibit some minor degree of surface defects, discovered during the annotation stage. The set of parts for defective cases was distributed randomly and evenly amongst the four systems. Defect instances were created either by manipulating manufacturing settings on the systems, in order to create manufacturing process-based defects such as bulges, gaps, warping, over extrusion and under extrusion. Surface defect examples of a mechanical nature, in contrast to those generated by process error, were generated by replicating the handling and manipulation errors. These examples result in a real manufacturing environment, including the generation of cuts, holes, cracks, marks, and scratches. The majority of instances of burrs and roughness defects were as a result of naturally occurring defective features on parts.
@@
Figure 2: Experimental setup in laboratory conditions, showing mounting of the MechMind Nano, Intel RealSense and Stereolabs Zed2i sensors on an ABB IRB 120 robot arm, oriented towards the mounting fixture. Not pictured: MechMind LSR-L camera, mounted out of view above the cell frame.
@@
Data was acquired in a real-world environment in a robotics development cell. (Figure 2). Within this cell, an ABB IRB 120 robot arm was used to mount and orient the MechMind-Nano, Intel RealSense and Stereolabs Zed2i sensors, while MechMind LSR-L sensor was fixed to the upper exterior of the cell enclosure. The positions of all four industrial 3D imaging sensors were static throughout the data acquisition process. This mounting configuration was used to ensure that all four sensors would be operating within their optimal scanning ranges (as described in Table I), and all four sensors were oriented towards a mounting fixture that held the parts in place during scanning. The mounting fixture allowed 2-axis rotation, with a 360-degree rotation around its primary axis and a 90-degree range of motion in its secondary axis. In each scanning procedure, the part was fixed at the centre of this fixture, and allowed to rotate through a range of orientations, with a minimum of a full rotation of the primary axis at 45-degree intervals, with the secondary axis fixed at both 0 and 30 degrees, respectively. This provided, in most instances, a complete view of the part. In those instances where additional coverage was desired, further variance in the secondary axis position was used to achieve the complete view.
@@
Each of the industrial scanners deployed captured six-channel images in the capture resolutions defined in Table I, with these channels representing the r, g, b, x, y and z coordinates with respect to the sensors’ local reference coordinate frame. These [x, y, z] vectors represent the point cloud scan obtained by the sensor, and in all instances are captured as .PLY files. Additionally, for all sensors, the corresponding [r, g, b] values for each pixel are captured as a PNG RGB image, which enables us to have a 1-to-1 mapping between the RGB and XYZ image data for each scan. The scene captured within our environment was open to indirect natural lighting sources during the period of scanning, as well as diffuse indirect artificial light sources when natural illumination levels were insufficient for proper data acquisition conditions. Thus, as scanning was conducted throughout the day, across weather and seasonal change variations, the data captured represents a wide range of natural and artificial lighting conditions as can be observed in Figure 1 and 3.
III-B Ground Truth Annotations